# Characterizations of voting rules based on majority margins

### v1

@@ 1 Introduction
Moreover, these distinctions are not merely academic. Examples from real political elections with ranked ballots are shown in Figure 2, where the standard version of Minimax selects Dornan, while the winning votes version selects Flanagan, and Figure 3, where the standard version selects Arab, while the winning votes version selects Worlobach; in each case, we only show candidates in the Smith set (Smith 1973), which is the smallest of candidates such that every candidate in the set has a positive margin against every candidate outside the set. Table 1 shows the frequencies with which the two versions of Minimax disagree in four election databases, conditional on their being no Condorcet winner (both versions of Minimax elect the Condorcet winner whenever one exists).
@@ 2.1 Preferential Equality
Recall from Section 1 that Instant Runoff Voting (IRV) violates Preferential Equality, at least hypothetically. In fact, this is far from hypothetical. Table 2 shows lower bounds [^6] The reason these frequencies are only lower bounds is that we did not exhaustively search for coalitions $I$ and $J$ as described in the text but rather used a heuristic to find such coalitions $I$ and $J$. on the frequencies in four election databases of profiles $\mathbf{P}$ that witness a violation by IRV of the coalitional version of Preferential Equality in the following sense: there exist candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ such that if, counterfactually, the voters in $I$ were to switch to ranking $y$ immediately above $x$, this would have a different effect than if instead the voters in $J$ were to make exactly the same switch. Table 2 also gives lower bounds on the frequencies of profiles $\mathbf{P}$ that witness a violation by IRV of the obvious coalitional version of Preferential Compensation, which is equivalent to Preferential Compensation: there exists candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ and $y$ immediately above $x$, respectively, such that each coalition flipping their ranking of $\{x,y\}$ changes the IRV outcome. All the frequencies are conditional on there not being a candidate who receives a majority of first-place votes in the first round, known as an absolute majority winner (AMW), since if there is an AMW, then IRV, Plurality voting, and Condorcet voting rules all agree. The frequencies are very high, showing that when IRV elections are interesting (i.e., there is no AMW), there are often coalitions with unequal power in terms of how they could have changed the election.
@@
Stable Voting Dataset 197 60 0.63 0.28
CIVS Dataset 1,883 590 0.61 0.33
PrefLib Politics (soi) 308 104 0.41 0.43
Otis 2022 dataset 450 194 0.44 0.39
Table 2: The relevant profiles are elections in which (i) every voter submits a linear order on some subset of the candidates, (ii) there is a unique IRV winner, and (iii) there are no ties when iteratively removing candidates with the fewest first-place votes to find the IRV winner. The last two columns contain lower bounds on the frequencies of relevant profiles witnessing a Preferential Equality Violation (PEV) or Preferential Compensation Violation (PCV) for IRV, conditional on there not being an absolute majority winner (AMW), in elections from the Stable Voting Dataset 12-20-2024 (Holliday and Pacuit 2024), CIVS Dataset 2024-12-15 (Myers 2024a, b), PrefLib Politics (soi files as of 12-20-2024) (Mattei and Walsh 2013), and Otis 2022. Note that we restricted to elections that are not “test” elections in the CIVS Dataset.
@@ 2.3 Proof of characterization
Suppose $\mathbf{P}$ is a profile satisfying the four conditions. By condition 1, we only need to show that $V(\mathbf{P})=V(\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}})$. For every $ab\in X^{(2)}$, let $\mathbf{P}_{|ab}$ be the profile [^7] This “profile” could be empty, but margins can be defined on the empty profile trivially. that keeps only the designated $ab$-voters in $\mathbf{P}$. Note that $\mathbf{P}$ is the disjoint union of $\mathbf{P}_{|ab}$ for all $ab\in X^{(2)}$. By conditions 1 and 2, for any $ab,cd\in X^{(2)}$ such that $\{a,b\}\neq\{c,d\}$, we have $\mathcal{M}_{\mathcal{P}_{cd}}(a,b)=0$, since $G_{cd}^{\top}$ and $G_{cd}^{\bot}$ rank $a,b$ differently. This means $\mathcal{M}_{\mathbf{P}}(a,b)=\mathcal{M}_{\mathbf{P}_{|ab}}(a,b)-\mathcal{M}_{\mathbf{P}_{|ba}}(a,b)=|V(\mathbf{P}_{|ab})|-|V(\mathbf{P}_{|ba})|$ for any $ab\in X^{(2)}$. By condition 4, at most one of $\mathbf{P}_{|ab}$ and $\mathbf{P}_{|ba}$ can be non-empty. This means that if $\mathcal{M}_{\mathbf{P}}(a,b)\leq 0$, then $V(\mathbf{P}_{|ab})$ must be empty, so trivially we have $V(\mathbf{P}_{|ab})=\{v(ab,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(a,b)/2,\star\in\{\top,\bot\}\}$. On the other hand, if $\mathcal{M}_{\mathbf{P}}(a,b)>0$, then $\mathbf{P}_{|ba}$ must be empty, meaning that $|V(\mathbf{P}_{|ab})|=\mathcal{M}_{\mathbf{P}}(a,b)$. By conditions 2 and 3, this means $V(\mathbf{P}_{|ab})$ must be $\{v(ab,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(a,b)/2,\star\in\{\top,\bot\}\}$. This completes the proof. ∎
@@ 4.3 Proof of characterization
Now let $\mathbf{Q}$ (resp. $\mathbf{Q}^{\prime}$) be the disjoint union of $\mathbf{P}$ (resp. $\mathbf{P}^{\prime}$) and all $\mathbf{A}_{D}$ (resp. $\mathbf{A}^{\prime}_{D}$) for $a,b\in X$ with $a\prec b$. Then #$\#_{\mathbf{Q}}=\#_{\mathbf{Q}^{\prime}}$. This is because, for any $x,y\in X$,
@@ Appendix A Appendix: Allowing Noncomparability
Let $\mathsf{LOSN}$ (resp. $\mathsf{WOSN}$) be the set of all profiles $\mathbf{R}$ such that for any $i\in V(\mathbf{R})$, $\mathbf{R}(i)$ is a LOSN (resp. WOSN) on $X(\mathbf{R})$. We call $S(\mathbf{R}(i))$ voter $i$’s side-noncomparability class.

### v2

@@ 1 Introduction
Moreover, these distinctions are not merely academic. Examples from real political elections with ranked ballots are shown in Figure 2, where the standard version of Minimax selects Dornan, while the winning votes version selects Flanagan, and Figure 3, where the standard version selects Arab, while the winning votes version selects Worlobach; in each case, we only show candidates in the Smith set (Smith 1973), which is the smallest set of candidates such that every candidate in the set has a positive margin against every candidate outside the set. Table 1 shows the frequencies with which the two versions of Minimax disagree in four election databases, conditional on their being no Condorcet winner (both versions of Minimax elect the Condorcet winner whenever one exists).
@@ 2.1 Preferential Equality
Recall from Section 1 that Instant Runoff Voting (IRV) violates Preferential Equality, at least hypothetically. In fact, this is far from hypothetical. Table 2 shows lower bounds [^6] The reason these frequencies are only lower bounds is that we did not exhaustively search for coalitions $I$ and $J$ as described in the text but rather used a heuristic to find such coalitions $I$ and $J$. on the frequencies in four election databases of profiles $\mathbf{P}$ that witness a violation by IRV of the coalitional version of Preferential Equality in the following sense: there exist candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ such that if, counterfactually, the voters in $I$ were to switch to ranking $y$ immediately above $x$, this would have a different effect than if instead the voters in $J$ were to make exactly the same switch. Table 2 also gives lower bounds on the frequencies of profiles $\mathbf{P}$ that witness a violation by IRV of the obvious coalitional version of Preferential Compensation, which is equivalent to Preferential Compensation: there exists candidates $x$ and $y$ and equally sized coalitions $I$ and $J$ of voters ranking $x$ immediately above $y$ and $y$ immediately above $x$, respectively, such that each coalition flipping their ranking of $\{x,y\}$ changes the IRV outcome. All the frequencies are conditional on there not being a candidate who receives a majority of first-place votes in the first round, known as an absolute majority winner (AMW), since if there is an AMW, then IRV, Plurality voting, and Condorcet voting rules all agree. The frequencies show that when IRV elections are interesting (i.e., there is no AMW), it is not unusual that there are coalitions with unequal power in terms of how they could have changed the election.
@@
Stable Voting Dataset 130 60 0.75 0.41
CIVS Dataset 1,216 590 0.66 0.57
PrefLib Politics (soi) 150 104 0.09 0.16
Otis 2022 dataset 448 194 0.17 0.19
Table 2: The relevant profiles are obtained from elections in which (i) every voter submits a linear order on a subset of the candidates, (ii) there is a unique IRV winner, and (iii) there are no ties when iteratively removing candidates with the fewest first-place votes; for each of these elections, we restrict the profile to the “top three” candidates, defined as the IRV winner plus the last two losers to be eliminated. The last two columns contain lower bounds on the frequencies of relevant profiles witnessing a Preferential Equality Violation (PEV) or Preferential Compensation Violation (PCV) for IRV, conditional on there not being an absolute majority winner (AMW), in elections from the Stable Voting Dataset 12-20-2024 (Holliday and Pacuit 2024), CIVS Dataset 2024-12-15 (Myers 2024a, b), PrefLib Politics (soi files as of 12-20-2024) (Mattei and Walsh 2013), and Otis 2022. Note that we restricted to elections that are not “test” elections in the CIVS Dataset.
@@ 2.3 Proof of characterization
Suppose $\mathbf{P}$ is a profile satisfying the four conditions. By condition 1, we only need to show that $V(\mathbf{P})=V(\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}})$. For every $ab\in X^{(2)}$, let $\mathbf{P}_{|ab}$ be the profile [^7] This “profile” could be empty, but margins can be defined on the empty profile trivially. that keeps only the designated $ab$-voters in $\mathbf{P}$. Note that $\mathbf{P}$ is the disjoint union of $\mathbf{P}_{|ab}$ for all $ab\in X^{(2)}$. By conditions 1 and 2, for any $ab,cd\in X^{(2)}$ such that $\{a,b\}\neq\{c,d\}$, we have $\mathcal{M}_{\mathbf{P}_{|cd}}(a,b)=0$, since $G_{cd}^{\top}$ and $G_{cd}^{\bot}$ rank $a,b$ differently. This means $\mathcal{M}_{\mathbf{P}}(a,b)=\mathcal{M}_{\mathbf{P}_{|ab}}(a,b)-\mathcal{M}_{\mathbf{P}_{|ba}}(a,b)=|V(\mathbf{P}_{|ab})|-|V(\mathbf{P}_{|ba})|$ for any $ab\in X^{(2)}$. By condition 4, at most one of $\mathbf{P}_{|ab}$ and $\mathbf{P}_{|ba}$ can be non-empty. This means that if $\mathcal{M}_{\mathbf{P}}(a,b)\leq 0$, then $V(\mathbf{P}_{|ab})$ must be empty, so trivially we have $V(\mathbf{P}_{|ab})=\{v(ab,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(a,b)/2,\star\in\{\top,\bot\}\}$. On the other hand, if $\mathcal{M}_{\mathbf{P}}(a,b)>0$, then $\mathbf{P}_{|ba}$ must be empty, meaning that $|V(\mathbf{P}_{|ab})|=\mathcal{M}_{\mathbf{P}}(a,b)$. By conditions 2 and 3, this means $V(\mathbf{P}_{|ab})$ must be $\{v(ab,\star,k)\mid 1\leq k\leq\mathcal{M}_{\mathbf{P}}(a,b)/2,\star\in\{\top,\bot\}\}$. This completes the proof. ∎
@@ 4.3 Proof of characterization
Note that since by assumption $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{P}^{\prime}}$, $d_{D}$ is also equal to #$\#_{\mathbf{P}^{\prime}}(b,a)-\#_{\mathbf{P}}(b,a)$. So, as long as $D=\{x,y\}$, $d_{D}=\#_{\mathbf{P}^{\prime}}(x,y)-\#_{\mathbf{P}}(x,y)$.
@@
Now let $\mathbf{Q}$ (resp. $\mathbf{Q}^{\prime}$) be the disjoint union of $\mathbf{P}$ (resp. $\mathbf{P}^{\prime}$) and $\mathbf{A}_{D}$ (resp. $\mathbf{A}^{\prime}_{D}$) for all doubletons $D\subseteq X$. Then #$\#_{\mathbf{Q}}=\#_{\mathbf{Q}^{\prime}}$. This is because, for any $x\not=y\in X$,
@@ Appendix A Appendix: Allowing Noncomparability
Let $\mathsf{LOSN}$ (resp. $\mathsf{WOSN}$) be the set of all profiles $\mathbf{R}$ such that for any $i\in V(\mathbf{R})$, $\mathbf{R}(i)$ is an LOSN (resp. WOSN) on $X(\mathbf{R})$. We call $S(\mathbf{R}(i))$ voter $i$’s side-noncomparability class.